OpenStudy (anonymous):

please show the steps to find the first derivative of P(t)=500/(1+e^(-t) )

OpenStudy (superdavesuper):

check ur eqn again plz? there is no x in ur expression for f(x)

OpenStudy (anonymous):

sorry its a population question so its P(t)

OpenStudy (superdavesuper):

OK that makes more sense :) u can rewrite it: P(t)=500*[1+e^(-t)]^-1 does it make the differentiation a lil easier?

OpenStudy (anonymous):

oh I already got that step im not sure what to do after :/ (500)'(1+e^-t)+(500)[(1+e^-t)] (1+e^-t)^-1+500[(-1(1+e^-t)^-2] Im confused after that

OpenStudy (superdavesuper):

OK substitute 1+e^-t = y ok? so P(t)=500*y^-1 right?

OpenStudy (superdavesuper):

then use dP/dt=dP/dy*dy/dt.....

OpenStudy (anonymous):

and then solve it and resubstitute e back?

OpenStudy (superdavesuper):

well yes but u need to use the eqn above too - u have seen it from ur book right?

OpenStudy (anonymous):

right ya

OpenStudy (superdavesuper):

so what u got? :)

OpenStudy (anonymous):

for derivating (1+e^-t)^-1 I used the power rule -1(1+e^-t)^-2 (-e^-t) [(1+e^-t)(-e^-t)]^-2

OpenStudy (superdavesuper):

u are jumping ahead, emily!! remember dP/dt=dP/dy*dy/dt P=500*y^-1 use Power rule to find what dP/dy is first plz?

OpenStudy (anonymous):

ok so to derive P= 500*y^-1 P'(x)= -1(500)y^-2?

OpenStudy (superdavesuper):

should be P'(y) but close enough :) okay we get dP/dy; now we get to dy/dt... y=1+e^-t so what is dy/dt?

OpenStudy (anonymous):

thats where im having difficulty :/ I dont know

OpenStudy (superdavesuper):

do u know d(e^x)/dx=e^x?

OpenStudy (anonymous):

oh i missed that lesson so i dnt quite understand it so is it like (1+e^-t)[(-1)(e^-t)]?

OpenStudy (superdavesuper):

read up about d(e^x)/dx=e^x from the link below: http://tutorial.math.lamar.edu/Classes/CalcI/DiffExpLogFcns.aspx

OpenStudy (superdavesuper):

dy/dt=d(1)/dt+d(e^-t)/dt = 0 + (-1)e^-t =-e^-t

OpenStudy (anonymous):

ok im trying my best :/ (1+e^-t)^-1+[500(-e^-t)^-2]

OpenStudy (superdavesuper):

almost there, emily....i promise ok?

OpenStudy (superdavesuper):

just refresh on what we have done: dP/dt = dP/dy * dy/dt dP/dy = -1(500)y^-2 dy/dt = -e^-t with me so far?

OpenStudy (anonymous):

ok lol um im giving up for now :/ but i do see that you have to mltiply them together -1(500)y^-2(-e^-t)?

OpenStudy (superdavesuper):

hahahaa why u giving up? dats ur ans right there, emily!!! :)

OpenStudy (anonymous):

oh thank god

OpenStudy (superdavesuper):

see its NOT that difficult :P